# MOS-Attack: A Scalable Multi-objective Adversarial Attack Framework

### v2

@@ 4.3 Implementation: Loss Functions
The final step of implementing our attack is to specify multiple surrogate loss functions. We incorporate a selection of significant loss functions that are well-documented in existing literature [34, 10, 46], along with innovative loss functions that have been identified through rigorous exploration in the domain of loss search [50, 30]. Details of these loss functions can be found in Table 1 and Table 2.
@@ 5.1 Experiment Setup
Datasets and Models. We employed 17 distinct from RobustBench [12], which includes 12 models [47, 37, 40, 36, 4, 27] trained on the CIFAR-10 [28] dataset and 5 models [49, 39] based on ImageNet [15] dataset. For performance evaluation, we used all 10,000 test images from the CIFAR-10 validation dataset and 5,000 images from ImageNet validation dataset. To enable direct comparison with the reported accuracy of the ACG attack, we preserved the same image indexing for the ImageNet dataset as [51].
@@
Attack Parameters. Notably, the number of iterations for our implemented attacks, including MOS Attack, APGD-CE, and APGD-All, are uniformly set to 50. This choice ensures thorough and rigorous testing of all methods. Additionally, the remaining attack parameters follow the same configuration as outlined in APGD [11].
@@ 5.2 Overall Results
ID Architecture MOS-8 (1) MOS-8 (8) Upper Bound Diff.
@@ 5.4.1 Loss Synergistic Pattern
Figure 2 depicts the synergistic loss patterns for CIFAR-10 and ImageNet. A significant portion of the adversarial examples—$61.3\%$ for CIFAR-10 and $57.8\%$ for ImageNet—contribute to all loss functions, indicating that the majority of solutions optimize them concurrently. This observation suggests a low level of conflict among the loss functions and helps explain why employing a single loss function ($K=1$) can yield near-optimal results.

### v3

@@ 4.3 Implementation: Loss Functions
The final step of implementing our attack is to specify multiple surrogate loss functions. We incorporate a selection of significant loss functions that are well-documented in existing literature [madry:2018:towards, croce:2020:minimally, wang:2020:improving], along with innovative loss functions that have been identified through rigorous exploration in the domain of loss search [xia:2024:tightening, li:2022:autoloss]. Details of these loss functions can be found in Table 1 and Table 2.
4.4 Runtime/Efficiency Comparison
Theoretically, MOS-Attack with $N$ loss functions and $N$ samples incurs only a constant factor more in computational expense compared to conducting $N$ single-objective attacks. Thus, the MOS-8(5) Attack may potentially offer greater efficiency than ALL-8(8) (5 v.s. 8 examples).
Table 4: Complexity analysis of gradient computation of single-objective and set-based optimization.
Method Gradient Computation Complexity
Single-objective $\frac{\partial L_{i}}{\partial{\bm{x}}_{i}}=\underbrace{\frac{\partial L_{i}}{\partial{\bm{h}}_{i}}}_{\begin{subarray}{c}\text{Vector}\\ \in\mathbb{R}^{m}\end{subarray}}\cdot\underbrace{\frac{\partial{\bm{h}}_{i}}{\partial{\bm{x}}_{i}}}_{\begin{subarray}{c}\text{Jacobian}\\ \in\mathbb{R}^{m\times d}\end{subarray}},\forall i$ $\mathcal{O}\left(N\cdot\text{Cost}\left(\frac{\partial{\bm{h}}}{\partial{\bm{x}}}\right)\right)$
Set-based $\frac{\partial g}{\partial\bm{X}}=\underbrace{\frac{\partial g}{\partial\bm{L}}}_{\begin{subarray}{c}\text{Vector}\\ \in\mathbb{R}^{1\times N}\end{subarray}}\cdot\underbrace{\frac{\partial\bm{L}}{\partial{\bm{h}}}}_{\begin{subarray}{c}\text{Matrix}\\ \in\mathbb{R}^{N\times m}\end{subarray}}\cdot\underbrace{\frac{\partial{\bm{h}}}{\partial\bm{X}}}_{\begin{subarray}{c}\text{Jacobian}\\ \in\mathbb{R}^{m\times d\times N}\end{subarray}}$ $\underbrace{\mathcal{O}\left(N\cdot\text{Cost}\left(\frac{\partial{\bm{h}}}{\partial{\bm{x}}}\right)\right)}_{\text{Gradient}}+\underbrace{\mathcal{O}\left(Nm\right)}_{\text{Matrix-Vector Multiplication}}$
The gradient computation cost of $\frac{\partial g}{\partial\bm{L}}$ and $\frac{\partial\bm{L}}{\partial{\bm{h}}}$ is negligible compared to $\frac{\partial{\bm{h}}}{\partial{\bm{x}}}$.
@@ 5.1 Experiment Setup
Datasets and Models. We employed 17 distinct from RobustBench [croce:2021:robustbench], which includes 12 models [wang:2023:better, rebuffi:2021:fixing, sehwag:2022:robust, rade:2022:reducing, augustin:2020:adversarial, huang:2021:exploring] trained on the CIFAR-10 [CIFAR10] dataset and 5 models [wong:2020:fast, salman:2020:do] based on ImageNet [deng:2009:imagenet] dataset. For performance evaluation, we used all 10,000 test images from the CIFAR-10 validation dataset and 5,000 images from ImageNet validation dataset. To enable direct comparison with the reported accuracy of the ACG attack, we preserved the same image indexing for the ImageNet dataset as [yamamura:2022:diversified].
@@
Attack Parameters. Notably, the number of iterations for our implemented attacks, including MOS Attack, APGD-CE, and APGD-All, are uniformly set to 50. This choice ensures thorough and rigorous testing of all methods. Additionally, the remaining attack parameters follow the same configuration as outlined in APGD [croce:2020:reliable].
@@ 5.2 Overall Results
ID Architecture
MOS-8
MOS-8
Upper
Bound
Diff.
@@ 5.4.1 Loss Synergistic Pattern
[…]
@@ 6 Conclusion
Acknowledgements
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China [GRF Project No. CityU11215622, CityU11212524], and by Key Basic Research Foundation of Shenzhen, China (JCYJ20220818100005011).